### v1

# Fishers Harvest Parallel Unlearning in Inherited Model Networks

Abstract.
Unlearning in various learning frameworks remains challenging, with the continuous growth and updates of models exhibiting complex inheritance relationships. This paper presents a novel unlearning framework, which enables fully parallel unlearning among models exhibiting inheritance. A key enabler is the new Unified Model Inheritance Graph (UMIG), which captures the inheritance using a Directed Acyclic Graph (DAG). Central to our framework is the new Fisher Inheritance Unlearning (FIUn) algorithm, which utilizes the Fisher Information Matrix (FIM) from initial unlearning models to pinpoint impacted parameters in inherited models. By employing FIM, the FIUn method breaks the sequential dependencies among the models, facilitating simultaneous unlearning and reducing computational overhead. We further design to merge disparate FIMs into a single matrix, synchronizing updates across inherited models. Experiments confirm the effectiveness of our unlearning framework. For single-class tasks, it achieves complete unlearning with 0% accuracy for unlearned labels while maintaining 94.53% accuracy for retained labels on average. For multi-class tasks, the accuracy is 1.07% for unlearned labels and 84.77% for retained labels on average. Our framework accelerates unlearning by 99% compared to alternative methods.
1. Introduction
The explosive growth of data volume and the increasing complexity of applications necessitate models that can constantly update to adapt to new data, constraints, and standards (Jayaraman and Evans, 2019). Accordingly, the need for data privacy, the auditing of illegal information from models, and compliance with regulations and industry standards arises. This leads to the concept of “model forgetting” or “unlearning,” which involves removing specific information from trained models.
The importance of unlearning lies in its ability to remove data that may raise privacy concerns, ensuring that the model can no longer recognize or rely on such data, in compliance with regulations such as the General Data Protection Regulation (GDPR) (Voigt and Von dem Bussche, 2017). Additionally, unlearning plays a crucial role in data correction, removing the impact of erroneous or biased data from the model and enhancing its accuracy and fairness. This not only protects user privacy but also boosts the model’s quality and reliability, aligning with both ethical standards and legal requirements.
Machine unlearning becomes particularly challenging when models are consistently recorded with a directed acyclic graph (DAG) topology, where models inheriting from predecessors also require updates. Examples of such scenarios include a basic model being customized to meet various Machine Learning Operations (MLOps) lifecycle requirements (Matsui and Goya, 2023; Kreuzberger et al., 2023) or distilled to fit computational resource constraints (Polino et al., 2018).
The concept of model inheritance can be applied to prevalent learning frameworks, such as Federated Learning (FL) (Li et al., 2020a), Distributed Data-Parallel Learning (DDPL) (Li et al., 2020c), Incremental Learning (IL) (Van de Ven et al., 2022), and Transfer Learning (TL) (Weiss et al., 2016). In this case, unlearning even a single model would need to unlearn the entire subgraph rooted at the origin to ensure all descendant models are consistent and relevant.
Research Gap. Various machine unlearning techniques have been proposed. Unfortunately, the inheritance relationships among related models prevent most existing unlearning methods from being executed in parallel, including re-training (Kim and Woo, 2022; Liu et al., 2024b), gradient ascent (Zhang et al., 2024), or knowledge distillation methods (Sinha et al., 2023). On the other hand, Sharding, Isolation, Slicing, and Aggregated Training (SISA) (Bourtoule et al., 2021) is a commonly used efficient method that reduces the dependence between data by dividing it into multiple segments and training multiple models in parallel. However, SISA is also inapplicable when there are inheritance relationships between models (Liu et al., 2024a). This is due to the fact that SISA needs to be performed sequentially from the starting point of updating to the end of the subgraph, incurring significant overhead and resource consumption during unlearning. In this sense, there is a pressing need for an effective and efficient way to perform unlearning in the presence of model inheritance relationships.
Research Question. To address this research gap, we need to identify the commonalities across various prevalent frameworks in unlearning and determine how a new unlearning design can be universally applied. Specifically, we raise two key Research Questions (RQs):
RQ1. How can unlearning requests be swiftly allocated to models that need updates and then prepare these models for unlearning tasks within a large-scale, interdependent model network?
RQ2. How can the model unlearning tasks be efficiently and effectively executed, especially when the unlearning of one model impacts subsequent models?
In response to these two RQs, we investigate the topological structure and unlearning impact of inherited models in four prevalent frameworks: FL, DDPL, IL, and TL. Specifically, we study the key factors affecting the efficiency, accuracy, and consistency of model unlearning. These include the selection of model parameters, changes in data distribution, and utilization of computational resources.
Contributions. This paper presents a novel parallel unlearning framework that can conduct efficient unlearning among large numbers of dependent models with complex inheritance relations, e.g., in FL, DDPL, IL, and TL. We first propose a new Unified Model Inheritance Graph (UMIG), to unveil the inheritance characteristic among models in various learning frameworks. With the assistance of the UMIG and Fisher Information Matrix (FIM), we develop the Fisher Inheritance Unlearning (FIUn) method that can efficiently pinpoint critical model parameters to be unlearned.
This allows independent unlearning adjustments on large-scale inherited models. To further enable effective one-shot unlearning of knowledge inherited from multiple upstream, we design a merging FIM function that consolidates disparate FIMs and identifies the critical parameters that need updating across all sources.
The contributions of this paper are summarized as follows:
  • We propose the UMIG, which abstracts various prevalent learning frameworks, e.g., FL, DDPL, IL, and TL, in a unified fashion using DAGs to represent models’ interdependence. This facilitates the analysis of large-scale inheritance-oriented model networks.
  • We design the FIUn method, which enables efficient parallel unlearning within the UMIG. The method leverages the FIM to quantify the significance of model parameters to unlearning tasks, and selectively scales the parameters according to their importance.
  • We further develop the merging FIM function to consolidate the FIMs from multiple upstream unlearning models into a cohesive matrix, thereby aggregating the unlearning tasks inherited from those models. This design facilitates the one-shot removal of inherited knowledge and reduces computational overhead within the UMIG.
Comprehensive experiments demonstrate that our method achieves complete unlearning for single-class unlearning tasks across all considered learning frameworks and model types, retaining an average accuracy of 94.53%. For multi-class unlearning tasks, the average unlearning accuracy drops to 1.07%, while the accuracy for retained labels reaches up to 84.77%. Additionally, our algorithm shows a 99% speed improvement across all considered frameworks compared to alternative methods.
Paper Organization. The rest of the paper is organized as follows. Section 2 reviews the preliminaries. Section 3 provides the proposed UMIG, followed by introducing the novel FIUn method and merging FIM function in Section 4. Section 5 presents the experimental settings. Section 6 presents our comprehensive experimental results based on four learning frameworks. Section 7 reviews the related works. Section 8 concludes this work.
2. Preliminaries
This section introduces the concepts, notation, and formulas of this paper, including machine unlearning and FIM.
2.1. Machine Unlearning
Machine unlearning refers to removing the influence of specific data subsets from a trained model while maintaining its performance on the remaining data (Foster et al., 2024a; Umer and Polikar, 2021). In other words, the goal is to eliminate the impact of specific labels on the model’s predictions while ensuring that the prediction accuracy for the remaining labels and the overall performance of the model remain unaffected.
Let $D=\left\{\left(x_{i},y_{i}\right)\right\}^{N}_{i=1}$ represent a dataset where $x_{i}$ is the $i$-th training sample and $y_{i}\in\left\{1,\cdots,K\right\}$ is the corresponding class label. Let $C_{f}$ denote the set of labels to be unlearned, corresponding to the data set $D^{f}\subset D$ to be unlearned, and $C_{r}$ represent the set of retained labels, corresponding to the remaining data set $D^{r}=D\backslash D^{f}$.
Unlearning aims to eliminate the impact of $D^{f}$ with labels in $C_{f}$ from the model while preserving its performance on $D^{r}$ with labels in $C_{r}$. In this context, we define $\phi_{\theta}\left(\cdot\right):X\rightarrow Y$ as a function parameterized by $\theta\in\mathbb{R}^{m}$, where $X\in\mathbb{R}^{n}$ and $Y\in\mathbb{R}^{K}$. The $k$-th component of $\phi_{\theta}\left(x\right)$ represents the probability that sample $x$ belongs to class $k$. The objective is to adjust $\theta$ such that the model effectively “forgets” the influence of $C_{f}$ while retaining accurate predictions on $C_{r}$.
2.2. Fisher Information Matrix
The FIM is a useful tool in statistics for assessing the accuracy of parameter estimates (Kirkpatrick et al., 2017; Golatkar et al., 2020). It quantifies the impact of parameter changes on the probability distribution of observed data.
Given a probability density function $p\left(x|\theta\right)$ conditioned on the model parameter $\theta$, an FIM $I\left(\theta\right)$ is the expected value of the second derivative of the negative log-likelihood function $\ell\left(\theta\right)$, as given by
(1) $I\left(\theta\right)=\mathbb{E}\left[-\frac{\partial^{2}\ell\left(\theta\right)}{\partial\theta^{2}}\right],\ \text{where}\ \ell\left(\theta\right)=\ln p\left(x|\theta\right).$
To simplify computation, in practice, the diagonal of the FIM can approximate the second derivative of the loss function (Steven, 1993; Aich, 2021). An equivalent form of $I\left(\theta\right)$ is given by
(2) $I(\theta)=\mathbb{E}\left[\left(\left(\frac{\partial\ell\left(\theta\right)}{\partial\theta}\right)\left(\frac{\partial\ell\left(\theta\right)}{\partial\theta}\right)^{\top}\right)\right]\ .$
In this paper, the FIM is employed to quantify the importance of data to the model parameters. By comparing the FIMs of the training dataset and the to-be-unlearned dataset, we can identify the critical parameters related to the knowledge that needs to be unlearned.
3. Capturing Model-Inheritance: UMIG
In this section, we explore the topological structure of four prevalent frameworks, FL, DDPL, IL, and TL, and analyze the impact of unlearning on the inherited models within these frameworks. We propose the UMIG, which can describe the model inheritance and update relationships in these learning frameworks in a unified manner.
(a) FL vanilla and DDPL
(b) Multi-layer/server FL
(c) IL and TL
(d) Decentralized FL
Figure 1. These diagrams demonstrate various learning frameworks that can be constituted to the proposed UMIG.
3.1. Diverse Learning Frameworks and Tasks
3.1.1. Federated Learning
A central server creates a global model, which participants download to locally update the model parameters based on their local data. The participants then send back their updated parameters. The central server aggregates the parameters to update and redistribute the global model. This process repeats until the model converges (McMahan et al., 2017; Li et al., 2020b; Krauß et al., 2024). As shown in Figure 1(a), the topology of FL consists of nodes representing model states and edges representing parameter transmission and aggregation. Updates to model $A$ trigger updates in the global model $D$ and then to $E$, $F$, $G$, and $H$. FL also has other structures.
Multi-layer FL. This variant introduces an intermediate aggregation layer to traditional FL (Chu et al., 2024; Li et al., 2021), as depicted in Figure 1(b). This approach adds multiple servers that perform intermediate aggregation between the clients and the global server, thus reducing the burden on the global server, improving system scalability, and enhancing model training.
DAG-FL. As shown in Figure 1(d), multiple clients update and aggregate models hierarchically in DAG-FL (Cao et al., 2021a, b; Yu et al., 2023). Each node can receive model updates from multiple parent nodes and send the updated results to multiple child nodes. This structure enables efficient model training and more flexible communication, better adapting to complex network environments and different application needs.
3.1.2. Distributed Data-Parallel Learning
Training tasks are typically segmented into multiple sub-tasks, distributed among various computing nodes or servers for execution, as documented in (Ge et al., 2023; Dean et al., 2012; Abadi et al., 2016; Paszke et al., 2019; Li et al., 2014). Each node independently processes its designated sub-task and periodically synchronizes its model parameters with others to collectively refine the global model. This iterative process is crucial, particularly in DDPL topologies — similar to the FL topology depicted in Figure 1(a) — which require the sub-tasks $E$, $F$, and $H$, as well as the aggregated model $I$, to be updated continually with each new iteration of the global model $D$.
3.1.3. Incremental Learning
Incremental (or “Continual”) learning entails training a model initially on a dataset and then progressively updating it as new data arrives, ensuring adaptability to the latest information (Han et al., 2023; Rajapaksha et al., 2023). IL can be visualized as shown in Figure 1(c), where nodes represent different updated states of the model, and edges depict the transition between model states after data updates. If the initial state model $G$ continues to update, then the subsequent models $A$ and $B$ that receive new data, as well as their related models, also need updates.
3.1.4. Transfer Learning
TL involves pre-training a base model on one dataset and then applying this model to a new task. Typically, this process includes fine-tuning the parameters of the last few layers to better suit the new task (Sendner et al., 2023; Rajapaksha et al., 2023). The model continues to train on the specific dataset of the new task to adapt fully to its requirements. Through ongoing fine-tuning and training, the model is optimized for the best performance on the new task. The TL topology, similar to the IL topology shown in Figure 1(c), also follows a comparable model updating process.
In these diverse learning frameworks, the inheritance relationship between models is pivotal. When a base model is updated or optimized, all corresponding child models derived from it must also be updated. To accurately capture this inheritance relationship, we employ the DAG topology as an effective abstraction method to capture the dependencies among models in these learning frameworks. The use of the DAG ensures that when a single model is updated, the entire subgraph—originating from this model and encompassing all related child models—is appropriately updated.
3.2. Unified Model Inheritance Graph (UMIG)
Figure 2. The darker blue nodes are all model nodes inheriting from the starting nodes $n_{s}$ and $n_{f}$ in lighter blue. A model node $n_{\ast}$ corresponds to a model $w_{\ast}$.
To uniformly handle and analyze the inheritance and update relationships of models across various learning frameworks, including FL, DDPL, IL, and TL, we propose a DAG-based UMIG. It provides a comprehensive framework to represent components such as models, data, and tasks and their interdependencies within these frameworks. It also facilitates a structured and intuitive understanding of these complex model networks. Figure 2 illustrates the ability of UMIG to depict the inheritance relationships between models.
As illustrated in Figure 2, we define a weighted directional graph $\mathcal{G}=(N,E)$. Herein, $N=\left\{n_{1},n_{2},\ldots,n_{m}\right\}$ represents the model nodes in the network, more specifically, a set of model parameters in a specific state or task. Moreover, $E=\left\{e_{01},e_{02},\ldots,e_{mm}\right\}$ represents the inheritance relationships between model nodes, encompassing parameter transfer and task inheritance. For instance, in FL and DDPL, the edges may represent the transfer of parameters from local to global models. By contrast, in TL, the edges signify the inheritance of tasks from a pre-trained model to a fine-tuned model.
Based on the source and state of the model, we further categorize the nodes into two types:
  • Discovery Nodes refer to models trained on data with an unlearned label set, enabling these nodes to adapt to the updated classification ability within the corresponding subgraph of the DAG. The root model nodes $n_{s}$ and $n_{f}$ in Figure 2 belong to this type.
  • Inherited Nodes comprise the remaining nodes in the subgraph, except the discovery nodes. Inherited nodes point to the discovery or other inherited nodes, and focus on refining classification performance. The model nodes $n_{h},n_{j},n_{l},n_{p},n_{g},n_{i},n_{n}$ belong to this type.
3.3. UMIG-Assisted Unlearning
With an emphasis on knowledge inheritance and unlearning tasks, UMIG allows a model to retain and utilize previously learned useful information when receiving new data, achieving an effective combination of old and new knowledge. UMIG also allows the model to remove specific data for continuous optimization and performance improvement.
The UMIG-assisted unlearning consists primarily of the following two processes:
  • Locating unlearning subgraphs. Based on the data that needs to be unlearned, a breadth-first search is conducted to identify the discovery nodes, which then spawn the subgraphs associated with the discovery nodes.
  • Updating models within unlearning subgraphs. All model nodes, including discovery and inherited nodes within the identified unlearning subgraphs, are updated to unlearn the parts of their models affected by the data.
By performing a breadth-first search of the entire model network, one or more discovery nodes are identified based on unlearning requests related to specific data. These discovery nodes serve as the roots for the unlearning subgraphs. As depicted in Figure 2, the unlearning subgraphs may overlap since inherited nodes may inherit from multiple discovery nodes. For example, the discovery nodes, $n_{s}$ and $n_{f}$, lead to two overlapping subgraphs with inherited nodes. Overlapping inherited nodes, such as $n_{h},n_{j},n_{l},n_{p}$, would require the elimination of knowledge from multiple discovery nodes when unlearning due to their connections to different unlearning subgraphs. This situation is known as a multi-root scenario. In contrast, inherited nodes within one unlearning subgraph, such as $n_{g},n_{i},n_{n}$ in a single-root scenario, only need to eliminate the impact from a single discovery node.
Inheritance is not just a simple mapping from the discovery nodes to the inherited nodes. Instead, it is achieved through multiple paths and varying depths. Each path represents a route of knowledge transfer via iteration, and the number of paths and differences in depth determine the complexity of the inheritance relationship.
A complex structure emerges when multiple paths are at varying depths from an inherited node to its corresponding discovery node(s), e.g., $n_{f},n_{g},n_{i}$ or $n_{s},n_{f},n_{h},n_{j}$. If unlearning operations are conducted simultaneously in terms of depth (in other words, nodes at the same depth from the root(s) are processed simultaneously and rely on sequential dependencies), nodes such as $n_{i}$ and $n_{j}$ may require multiple updates. This situation is termed an imbalanced-path scenario, contrasting with a balanced-path scenario (e.g., $n_{f},n_{h},n_{i},n_{l}$), where each node requires only a single update due to uniform path lengths and synchronous processing at each depth. In cases where unlearning is conducted asynchronously, regardless of depth, both scenarios may require multiple updates.
Answer to RQ1: We propose the UMIG to effectively visualize model interdependence and capture the model inheritance in a structured way, leveraging the DAG structure. Assisted by UMIG, the models inheriting the knowledge to be removed can be rapidly identified. Moreover, the UMIG offers a systematic way to collect and prepare inheritance information, thus facilitating the unlearning process of inherited models. We also examine various learning frameworks of inherited models based on the numbers of their upstream models and the paths of knowledge inheritance to elucidate the emerging challenges and considerations of the unlearning process in inherited nodes due to model inheritance.
Figure 3. The unlearning process in $\mathcal{G}$ consists of five steps: 1) Preparation. The discovery nodes $\{n^{\scriptscriptstyle D}_{i}\}$ are identified by performing a breadth-first search, with unlearning data $\hat{D}^{\scriptscriptstyle D}_{i}$ containing one or more labels. Subsequently, the unlearning graph $\mathcal{\hat{G}}$ is located which root at $\{n^{\scriptscriptstyle D}_{i}\}$. 2) Calculate unlearning the unlearning FIMs $\{\hat{F}_{i}^{\scriptscriptstyle D}\}$ of $\{n^{\scriptscriptstyle D}_{i}\}$. 3) Calculate the model FIM $F_{j}$ of node $n_{j}$ in $\mathcal{\hat{G}}$. 4) Merge unlearning FIMs $\hat{F}^{\scriptscriptstyle D}_{j,k}$ for $n_{j}$ in the case of multiple discovery nodes according to the topology. 5) Update $n_{j}$ by comparing the merged unlearning FIM $\hat{F}_{j}^{\scriptscriptstyle M}$ and $F_{j}$. Steps 2, 3, 4 and 5 apply to all nodes in $\mathcal{\hat{G}}$
4. FIUn-based Parallel Unlearning
This section introduces a novel parallel unlearning method, Fisher Inheritance Unlearning (FIUn), designed to address the challenges associated with machine unlearning in large-scale UMIG model networks. A trusted learning network is considered where all learning contributors follow the designed unlearning process. The FIUn method operates under the general expression provided by the UMIG and leverages FIMs. It initiates the unlearning process by identifying specific model parameters that need updates, making the process more systematic. Derived from models in the unlearning graph, the training data, and the requested unlearning data, the FIMs are crucial in pinpointing essential model parameters and recommending their updated values.
Notably, the FIUn method allows for the independent application of unlearning steps to each impacted model, facilitating a swift and effective parallel unlearning across extensive model networks. The FIUn method is depicted in Figure 3, and the procedural details are given in Algorithm 1.
1
2Input: Model network $\mathcal{G}=(N,E)$, Unlearing task $\mathcal{T}$
3Parameter: $\gamma$, $\tau$, $\eta$
4Use breadth-first search to identify discovery nodes $\{n^{\scriptscriptstyle D}_{i}\}$ and consequential unlearning graph $\mathcal{\hat{G}}$
5 for $n^{\scriptscriptstyle D}_{i}$ in $\{n^{\scriptscriptstyle D}_{i}\}$ parallel do
6       Calculate unlearning FIM $\hat{F}^{\scriptscriptstyle D}_{i}$ with Eq. (3)
7 end for
8
9for $n_{j}$ in $\mathcal{\hat{G}}$ parallel do
10       Calculate model FIM $F_{j}$ with Eq. (4)
11       Merge unlearning FIMs $\{\hat{F}^{\scriptscriptstyle D}_{j,k}\}$ with Eq. (5)
12       Update model parameters of node $n_{j}$ with Eq. (6)
13 end for
14
Algorithm 1 Fisher Inheritance Unlearning (FIUn)
4.1. Preparation
Upon receiving an unlearning task $\mathcal{T}$, accompanied by the unlearning dataset $\hat{D}$, FIUn initiates by conducting a breadth-first search from the starting node on the model network $\mathcal{G}$ to identify the discovery nodes $\{n^{\scriptscriptstyle D}_{i}\}$. The discovery nodes are the first nodes exhibiting significant recognition of the unlearning data, with $n^{\scriptscriptstyle D}_{i}$ representing the $i$-th discovery node. Subsequently, FIUn delineates the unlearning graph $\mathcal{\hat{G}}$, which is rooted at the identified discovery nodes. FIUn then updates all nodes in the unlearning graph to unlearn the requested unlearning knowledge.
4.2. FIUn
The main idea of FIUn is to identify critical model parameters by comparing the FIM from the unlearning dataset with the FIM from the training dataset. The differences between these matrices highlight the necessary adjustments in model parameters to achieve effective unlearning. The FIUn process unfolds as follows.
4.2.1. Calculate unlearning FIMs of discovery nodes
The FIUn method initiates by assessing the impact of the unlearning data on the model parameters of the discovery nodes, utilizing the FIM as specified in Eq. (3). This assessment is conducted using the model parameters and the unlearning data (see line 5, Algorithm 1).
The unlearning FIM for the $i$-th identified discovery node, $n^{\scriptscriptstyle D}_{i}$, is denoted by $\hat{F}^{\scriptscriptstyle D}_{i}$ and is calculated as
(3) $\hat{F}^{\scriptscriptstyle D}_{i}\!\!=\!\mathbb{E}\!\left[\!\left(\!\!\left(\frac{\partial\ln p(\hat{D}_{i}^{\scriptscriptstyle D}|w^{\scriptscriptstyle L}_{i})}{\partial w^{\scriptscriptstyle L}_{i}}\right)\!\!\!\left(\frac{\partial\ln p(\hat{D}_{i}^{\scriptscriptstyle D}|w^{\scriptscriptstyle L}_{i})}{\partial w^{\scriptscriptstyle L}_{i}}\!\right)^{\!\!\!\intercal}\right)\!\!\Bigg{|}_{w^{\ast}_{\hat{D}_{i}^{\scriptscriptstyle D}}}\!\right],$
where $\hat{D}^{\scriptscriptstyle D}_{i}$ represents the unlearning data for $n^{\scriptscriptstyle D}_{i}$, $w^{\scriptscriptstyle L}_{i}$ denotes the last layer parameters of the model $w_{i}$ at node $n^{\scriptscriptstyle D}_{i}$, and $w^{\ast}_{\hat{D}_{i}^{\scriptscriptstyle D}}$ refers to the optimal parameters learned on $\hat{D}_{i}^{\scriptscriptstyle D}$.
The unlearning FIM highlights the importance of each parameter in the last layer with respect to the specific unlearning task (Lee et al., 2023; Zhang et al., 2021). In FIUn, the FIM calculation is restricted to the last layer parameters while other layers remain frozen, aimed at reducing the computational complexity associated with FIM calculations, which is a challenge in traditional FIM-based unlearning methods. Using the last layer for unlearning can focus on the most relevant features. Since the last layer directly affects the final decision of the model, adjusting its parameters can effectively facilitate unlearning goals while maintaining stable performance on data not subject to unlearning (Kim et al., 2024).
4.2.2. Calculate model FIMs of all nodes in unlearning graph
The FIUn method proceeds by calculating the model FIMs of all models within the unlearning graph to identify the impact of model parameters in each node (see line 8, Algorithm 1).
For the $j$-th node $n_{j}$ in the unlearning graph $\mathcal{\hat{G}}$, either a discovery node or an inherited node, the model FIM, denoted by $F_{j}$, is given by
(4) $F_{j}\!=\!\mathbb{E}\!\left[\!\left(\!\!\left(\frac{\partial\ln p(D_{j}|w^{\scriptscriptstyle L}_{j})}{\partial w^{\scriptscriptstyle L}_{j}}\right)\!\!\!\left(\frac{\partial\ln p(D_{j}|w^{\scriptscriptstyle L}_{j})}{\partial w^{\scriptscriptstyle L}_{j}}\right)^{\!\!\!\intercal}\right)\!\Bigg{|}_{w^{\ast}_{D_{j}}}\right]\!,$
where $D_{j}$ is the training data for $n_{j}$, $w^{\scriptscriptstyle L}_{j}$ denotes the last layer parameters of the model $w_{j}$ at node $n_{j}$, and $w^{\ast}_{D_{j}}$ denotes the optimal parameters learned on $D_{j}$.
The model FIM quantifies the knowledge about the training dataset embedded in the model parameters and highlights the significance of each parameter. Similar to the unlearning FIM, only the FIM for the last layer parameter is calculated, aiming to minimize computational complexity.
4.2.3. Merge unlearning FIMs for all nodes in the unlearning graph
The FIUn method estimates the unlearning FIM for all nodes within the unlearning graph, including both discovery and inherited nodes (see line 9, Algorithm 1). The unlearning FIMs from discovery nodes are merged to form a collective unlearning FIM for each node in the unlearning graph. This merger is crucial as the linkages between the unlearning data and parameters are embedded within the model and can persist through model inheritance (Foster et al., 2024b), especially in multi-root scenarios where an inherited node may trace back to several discovery nodes. A novel merging FIM function is introduced, denoted by $\Phi\left(\cdot\right)$, and is defined as
(5) $\hat{F}^{\scriptscriptstyle M}_{j}=\Phi\left(\{\hat{F}^{\scriptscriptstyle D}_{j,k}\}\right),$
where $\hat{F}^{\scriptscriptstyle M}_{j}$ is the merged unlearning FIM for the $j$-th node in the unlearning graph, the set $\{\hat{F}^{\scriptscriptstyle D}_{j,k}\}$ collects the unlearning FIMs of discovery nodes within $\mathcal{\hat{G}}$ that can be traced back from node $n_{j}$, and $k$ denotes the $k$-th discovery node reachable from $n_{j}$. In this paper, the element-wise maximum is employed for the merging FIM function. Specifically, $\hat{F}^{\scriptscriptstyle M}_{j}$ assembles the largest elements among the unlearning FIMs of discovery nodes in multi-root scenarios and takes the only unlearning FIM in single-root scenarios.
Figure 4. There are three discovery nodes that need to undergo the unlearning process, namely model $w_{j}$, model $w_{s}$ and model $w_{f}$. The labels to be unlearned are label 1 and label 2. However, model $w_{j}$ and model $w_{s}$ only have label 1, while model $w_{f}$ only has label 2. Model $w_{h}$ inherits from these three discovery nodes, so we merge the unlearning FIMs of the discovery nodes and perform the unlearning of label 1 and label 2 on model $w_{h}$ together.
As shown in Figure 4, the model $w_{h}$ inherits from models $w_{j}$, $w_{s}$, and $w_{f}$ and needs to remove all requested unlearning data, though each discovery node contains only a portion of the unlearning data. The FIM of each model reflects the sensitivity of its parameters to the unlearning data, which can be interpreted as parameter importance. By taking the maximum of the corresponding elements across these unlearning FIMs, the merged FIM encompasses all unlearning tasks from the discovery nodes and effectively identifies the parameters that need updating to achieve complete unlearning.
4.2.4. Update models in unlearning graph
In the final step, the FIUn method identifies which parameters need to be updated by comparing the merged FIM with the model FIM, as described in line 10 of Algorithm 1. The outcomes of this comparison are also utilized to scale the parameters. For node $n_{j}$ in the unlearning graph, the parameters are updated as
(6) $\hat{w}^{\scriptscriptstyle L}_{j,l}=\begin{cases}\min(\tau\frac{F_{j,l}}{\hat{F}^{\scriptscriptstyle M}_{j,l}},\eta)w^{\scriptscriptstyle L}_{j,l},&\text{if }\frac{\hat{F}^{\scriptscriptstyle M}_{j,l}}{F_{j,l}}>\gamma;\\ w^{\scriptscriptstyle L}_{j,l},&\text{if}\ \frac{\hat{F}^{\scriptscriptstyle M}_{j,l}}{F_{j,l}}\leq\gamma,\end{cases}$
where $w^{\scriptscriptstyle L}_{j,l}$ and $\hat{w}^{\scriptscriptstyle L}_{j,l}$ are the $l$-th parameter in the last layer before and after the unlearning update, respectively. Similarly, $F_{j,l}$ and $\hat{F}^{\scriptscriptstyle M}_{j,l}$ are the $l$-th element of the model FIM and merged unlearning FIM of node $n_{j}$, respectively. The hyperparameters, including $\tau$, $\eta$, and $\gamma$, balance the unlearning impact on the unlearning data and the accuracy of the retained data.
During the model updating phase, only the parameters of the last layer are updated while the other layers remain frozen, consistent with our design of the unlearning FIM and model FIM, where only the FIM of the last layer is calculated. Elements within the FIM indicate the significance of each parameter relative to the dataset. The threshold $\gamma$ identifies the critical parameters that are significant for the unlearning task, and these parameters are scaled down accordingly. The hyperparameter $\eta$ ensures minimal change to achieve the desired unlearning effect. Non-critical parameters are left unchanged to maintain accuracy on data that does not require unlearning.
4.3. Discussion
From the perspective of the FIM, the impact of data on specific model parameters is consistently localized at precise positions within the parameter space (Foster et al., 2024b). The stable localization of the FIM across iterations helps precisely locate the parameter locations for unlearning at the class, client, and sample levels using data tied to specific label sets. By fine-tuning $\tau$, $\eta$, and $\gamma$ in Eq. (6) to match the sample size of the class being unlearned, the task transits flexibly between the class-level (Foster et al., 2024b) and sample-level unlearning (Golatkar et al., 2020). Specifically, when the unlearning FIM mirrors the model’s FIM with a class-level setting, it becomes a client-level unlearning task. All types of unlearning are performed irrespective of the succession of inheritance relationships. Consequently, employing FIUn disrupts the sequential dependencies traditionally required for unlearning, thus enabling fully parallel unlearning processes across any number of subgraphs.
Not only can models across different subgraphs undergo parallel unlearning, but those within the same subgraph at various depths can also be unlearned in parallel, thanks to a feature referred to as Hyper-Distance in this paper. This Hyper-Distance is achieved without compromising the integrity of the inheritance relationships. Hyper-Distance ensures that both unlearned and retained knowledge is effectively managed along the inheritance paths. This efficiency is possible because the FIUn method utilizes FIMs for unlearning, enabling a class-wise inheritance approach. Each model effectively incorporates a union of label sets from its predecessors, ensuring that the inheritance is preserved. Simultaneously, by focusing on label differences, we achieve fully parallel unlearning—making depths and paths irrelevant to the process.
Answer to RQ2: The FIUn method enables independently removing knowledge from each inherited model. This method utilizes the FIM to obtain the Hyper-Distance property which breaks sequential dependencies among models, enabling fully parallel unlearning on inherited models and significantly decreasing the execution time of unlearning tasks. To address the complexity arising from multiple upstream models, the merging FIM function is developed to aggregate unlearning FIMs from those models, facilitating the efficient, one-shot removal of inherited knowledge.
We also map the FIUn method and merging FIM function to the UMIG-assisted unlearning scenarios described in Section 3.3. This demonstrates how the proposed approach efficiently facilitates unlearning through the property of Hyper-Distance. Converting Imbalance into Balance. This scenario is particularly significant due to the fact that existing unlearning methods suffer from inevitable sequential dependencies. FIUn leverages the Hyper-Distance property to circumvent the influences of depths and paths within the inheritance topology. Consequently, imbalanced-path scenarios can be managed as effectively as balanced-path scenarios, where only one update operation is needed, regardless of whether unlearning is conducted synchronously or asynchronously, i.e., only one unlearning FIM needs to be computed for each inherited node.
Converting Multi-root into Single-root. The merging FIM function is proposed to refine multi-root scenarios. This function merges the FIMs of the corresponding unlearned label sets from multiple discovery nodes into a single FIM, as shown in Figure 4. Specifically, overlapping inherited nodes require only one update operation to eliminate the impact of multiple discovery nodes by conducting a parameter-wise merger, where only the most impactful parameters take effect in this FIM. This function eliminates the need for multiple comparisons among various unlearning FIMs in multi-root scenarios, hence enhancing unlearning efficiency.
4.4. Implementation
4.4.1. System Setting
In practice, FIUn can be handled differently depending on the system architecture, as described in the following:
  • Centralized: Model FIMs are computed and attached along with the model being published and sent to the central coordinator. When an unlearning request is initiated, the central server performs unlearning by merging the unlearning FIMs of the discovery nodes and executes unlearning in parallel.
  • Decentralized: The discovery nodes publicly share Each unlearning FIM. Owners of each inherited node can then combine these unlearning FIMs with their own local models at will, according to specific unlearning requirements. This decentralized approach enables efficient and effective unlearning.
4.4.2. Complexity
The computational complexity of an unlearning task comprises the complexities of Steps 1, 3, and 4 (i.e., calculate unlearning FIM, merge unlearning FIMs and update models). The complexity of Step 2 is not considered, as the computation of the model FIM can be performed beforehand. The overall complexity for the unlearning task is $\mathcal{O}(|\hat{D}||L|+|\{n^{\scriptscriptstyle D}_{i}\}|+|L|)$. Breaking down the complexities, Eq. (3) and Eq. (4) update $|L|$ parameters per data entry, with complexities $\mathcal{O}(|\hat{D}^{\scriptscriptstyle D}_{i}||L|)$ and $\mathcal{O}(|D_{j}||L|)$, respectively, where $|\hat{D}^{\scriptscriptstyle D}_{i}|\leq|\hat{D}|$ and $|L|$ is the number of parameters in the last layer. The term $\mathcal{O}(|D_{j}||L|)$ is omitted as model FIMs can be calculated along with the model training prior to any unlearning tasks. Eq. (5) includes a comparison and merger of $K$ unlearning FIMs, leading to a complexity of $\mathcal{O}(K)$, where $K$ is the number of traceable discovery nodes in the unlearning graph $\mathcal{\hat{G}}$ from node $n_{j}$ and $K\leq|\{n^{\scriptscriptstyle D}_{i}\}|$. Eq. (6) updates $|L|$ model parameters, leading to a complexity of $\mathcal{O}(|L|)$. As the computation of the unlearning FIM for discovery nodes and the model updates can proceed in parallel, the overall complexity of the unlearning tasks is $\mathcal{O}(|\hat{D}||L|+|\{n^{\scriptscriptstyle D}_{i}\}|+|L|)$.
5. Experimental Settings
This section first presents the datasets used in the experiments, the models selected, and the performance metrics for accuracy evaluation, followed by the topology setups and the benchmarks for unlearning to evaluate the FIUn method.
Table 1. Baseline of Federated Unlearning Performance On CIFAR-100 Unlearning Speed Comparison
Model #$C_{f}$ Cumulative Unlearning Time (s)
Gradient ascent Finetune Distill Re-training FIUn
$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
AlexNet 1 0.97 1.05 1.26 0.54 0.75 1.08 82.5 153.65 294.32 29.70 59.06 89.62 0.09 0.11 0.13
2 0.99 1.17 1.39 0.57 0.77 1.15 80.49 155.63 315.52 28.69 57.81 84.97 0.11 0.16 0.13
10 1.1 1.27 1.68 0.53 1.95 2.31 75.3 147.6 326.8 21.96 42.37 63.82 0.11 0.14 0.15
ResNet18 1 2.45 2.89 3.42 0.74 1.32 1.74 70.31 129.44 247.6 30.36 61.40 90.94 0.68 1.00 0.99
2 2.02 2.2 2.78 0.76 1.22 1.79 76.55 139.55 259.53 28.30 57.39 84.06 0.69 1.03 1.09
10 2.02 2.58 3.67 0.85 1.54 1.84 75.12 150.35 267.34 22.03 44.83 66.30 0.76 1.14 1.04
(a) FL in ResNet18 of #$C_{f}=1$
(b) FL in ResNet18 of #$C_{f}=10$
(c) FL in AlexNet of #$C_{f}=1$
(d) FL in AlexNet of #$C_{f}=10$
Figure 5. Baseline of Federated Unlearning Performance On CIFAR-100.
5.1. Datasets and Models
Datasets. The experiment uses two datasets to evaluate the proposed FIUn method. These two datasets are benchmarks for image classification tasks, encompassing diverse image classification tasks.
  • CIFAR100 (Krizhevsky et al., 2009). This dataset consists of 60,000 color images in 100 classes, with 600 images per class. The images are of size $32\times 32$ pixels. The dataset is divided into 50,000 training images and 10,000 test images. Each class in the CIFAR-100 dataset belongs to one of 20 superclasses. There is a coarse label (superclass) and a fine label (specific class) per image.
  • TinyImageNet (Le and Yang, 2015). This dataset contains 200 classes, each with 500 training images, 50 validation images, and 50 test images, with a total of up to 100,000 images. The images are resized to $64\times 64$ pixels, making them smaller than those in the original ImageNet dataset but larger than those in CIFAR-100.
Models. We use the AlexNet model (Iandola et al., 2016), ResNet18 model (Odusami et al., 2021) and DenseNet161 model (Talo, 2019) to evaluate the impact of our proposed FIUn method on the accuracy of the unlearned label set $C_{f}$ and the retained label set $C_{r}$.
Performance Metrics. We use the training datasets to check the accuracy of the model to evaluate its practicality in experiments. This is a common and reasonable approach since it directly evaluates if the removed information still influences the model, as widely accepted in the field (Tarun et al., 2023; Kim et al., 2024; Golatkar et al., 2020).
  • Accuracy on unlearned labels ($AD_{f}$). The accuracy of the unlearned label set in the unlearning model ideally would be close to zero. We use the training datasets to test the effectiveness of unlearning.
  • Accuracy on retained labels ($AD_{r}$). The accuracy of a retained label set of the unlearning model. It is expected to be close to the accuracy of the original model before unlearning.
  • Cumulative unlearning time. The cumulative time required for each model to unlearn labels during the training process in machine learning and deep learning.
  • Accuracy difference index ($\bigtriangleup_{acc}$). The difference between $AD_{r}$ and $AD_{f}$ also serves as a direct metric to evaluate unlearning performance. The greater the difference, the more effective the unlearning is, with a maximum 1.
Table 2. Federated Unlearning Performance On TinyImageNet
Model #$C_{f}$ Metrics Original Model (%) Re-training Model (%) FIUn (%) Cumulative Unlearning Time (s)
Re-training FIUn
$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
DenseNet161 1 ${AD}_{r}\uparrow$ 99.99 99.99 99.81 99.99 99.99 98.80 88.03 77.64 76.02 999.41 1996.43 2997.64 2.51 4.59 4.95
${AD}_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00
4 ${AD}_{r}\uparrow$ 99.99 99.99 99.80 99.99 99.99 99.51 83.01 71.43 71.00 1004.64 2006.31 3006.14 3.16 4.69 4.89
${AD}_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00
6 ${AD}_{r}\uparrow$ 99.99 99.99 99.96 99.99 99.99 99.16 80.00 67.67 66.58 1019.43 2020.16 3018.64 3.10 4.96 4.75
${AD}_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.00 0.00 0.00
10 ${AD}_{r}\uparrow$ 99.99 99.99 99.72 99.71 99.99 99.99 71.68 58.01 55.80 1031.43 2029.64 3030.64 2.95 4.85 4.89
${AD}_{f}\downarrow$ 99.99 99.99 98.82 0.00 0.00 0.00 0.00 0.00 0.00
ResNet18 1 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.73 99.97 99.46 112.95 216.39 314.29 1.35 3.68 3.42
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
4 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.17 99.98 99.98 109.89 209.31 310.42 1.63 3.26 3.57
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
6 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 99.54 99.98 99.98 107.39 208.12 305.32 1.46 3.37 3.51
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
10 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.98 99.99 99.99 104.21 201.36 302.74 1.52 3.21 3.36
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
Table 3. Incremental Unlearning Performance On TinyImageNet
Model #$C_{f}$ Metrics Original Model (%) Re-training Model (%) FIUn (%) Cumulative Unlearning Time (s)
Re-training FIUn
$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
DenseNet161 1 ${AD}_{r}\uparrow$ 99.99 99.99 99.81 99.99 99.99 98.80 89.76 78.50 73.87 991.43 1995.36 3989.47 2.75 4.85 4.43
${AD}_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00
4 ${AD}_{r}\uparrow$ 99.99 99.99 99.80 99.99 99.99 99.51 83.98 65.87 59.11 1006.32 2004.75 4009.41 2.64 4.43 4.64
${AD}_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00
6 ${AD}_{r}\uparrow$ 99.99 99.99 99.96 99.99 99.99 99.16 86.81 61.02 55.04 1019.43 2019.50 4019.46 2.36 4.47 4.74
${AD}_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.00 0.00 0.00
10 ${AD}_{r}\uparrow$ 99.99 99.99 99.72 99.99 99.99 99.99 85.95 54.61 45.46 1035.36 2041.78 4039.75 2.65 4.46 4.75
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
ResNet18 1 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 98.66 96.33 96.07 120.32 221.75 439.85 1.37 3.75 3.43
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
4 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 98.84 96.34 96.70 115.32 217.33 435.17 1.41 3.34 3.64
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
6 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 98.82 99.99 98.59 110.32 211.35 431.37 1.47 3.46 3.75
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
10 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.33 97.51 98.45 108.32 207.43 428.43 1.32 3.56 3.57
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
5.2. Learning Framework Setup and Benchmarks
Figure 6. Models topologies considered in the experiments.
Targeted Frameworks. We detailed setups for four learning frameworks under the UMIG with varying dataset and label distribution conditions.
  • Federated Unlearning. This process uses the structure shown in Figure 6. The CIFAR100 and TinyImageNet datasets are each randomly divided into five parts. When the model $w_{g}$ contains labels that need to be unlearned, the models that inherit from it need to perform the unlearning operation. We select models $w_{g}$, $w_{a}$, and $w_{b}$ for experimental demonstration.
  • Distributed Data-Parallel Unlearning. This process uses the structure shown in Figure 6. The CIFAR100 and TinyImageNet datasets are each randomly divided into three parts, each executing a sub-task in parallel. When the model $w_{g}$ contains labels that need to be unlearned, each sub-task performs the unlearning operation in parallel. We select models $w_{g}$, $w_{a}$, and $w_{b}$ for experimental demonstration.
  • Incremental Unlearning. This process uses the structure shown in Figure 6. In the CIFAR100 dataset, the first model is trained on data with labels 0-90, and each inheriting model’s training dataset progressively includes two additional classes. In the TinyImageNet dataset, the first model is trained on data with labels 0-190, and similarly, each inheriting model’s training dataset progressively includes two additional classes. We select models $w_{g}$, $w_{a}$, and $w_{b}$ for experimental demonstration.
  • Transfer Unlearning. Since TL involves fine-tuning the last layer parameters, adjustments are focused on the last layer of the model. Experiments, as illustrated in Figure 6, involve the CIFAR100 dataset where a model $w_{g}$ is trained on labels 0-90 and then transferred to models $w_{a}$ and $w_{b}$ trained on labels 0-92 and 0-98, respectively. Similarly, for the TinyImageNet dataset, model $w_{g}$ is trained on labels 0-190 and then transferred to models $w_{a}$ and $w_{b}$, which are trained on labels 0-192 and 0-198, respectively.
Benchmark. The experiment specifically focuses on class-level unlearning tasks for clarity of presentation. We re-train the models that include unlearned labels and subsequent models influenced by these labels due to inheritance. We remove the data with unlearned labels from each model’s dataset. The remaining data is then used to individually train each model. The models are trained sequentially according to their inheritance relationships.
Hyperparameters. The hyperparameters used in the experiment are shown in Appendix .1. The table lists the learning rates, optimizers, and other hyperparameters for different models, as well as the hyperparameters of our proposed FIUn method. Throughout the experimental process, all the hyperparameters remain unchanged to ensure a fair comparison of our method among different models.
Table 4. Transfer Unlearning Performance On CIFAR-100
Model #$C_{f}$ Metrics Original Model (%) Re-training Model (%) FIUn (%) Cumulative Unlearning Time (s)
Re-training FIUn
$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
AlexNet 1 ${AD}_{r}\uparrow$ 97.16 95.61 92.20 98.17 95.98 91.32 91.40 86.24 83.99 20.31 40.49 42.36 0.09 0.11 0.11
${AD}_{f}\downarrow$ 96.66 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
2 ${AD}_{r}\uparrow$ 97.16 95.61 92.20 96.88 96.68 93.16 82.36 80.47 79.99 21.20 40.56 41.92 0.08 0.13 0.11
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
4 ${AD}_{r}\uparrow$ 97.16 95.61 92.20 94.67 96.45 93.45 71.59 71.43 76.06 20.30 39.76 40.60 0.09 0.11 0.11
${AD}_{f}\downarrow$ 96.61 98.30 98.30 0.00 0.00 0.00 0.00 0.00 0.00
10 ${AD}_{r}\uparrow$ 97.16 95.61 92.20 99.99 99.99 99.99 75.97 69.56 76.15 19.42 39.46 39.35 0.08 0.13 0.11
${AD}_{f}\downarrow$ 98.57 99.28 99.28 0.00 0.00 0.00 20.28 20.28 20.28
ResNet18 1 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 89.09 93.51 95.22 22.21 41.52 42.93 0.15 0.24 0.24
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
2 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 85.89 94.89 95.79 21.19 41.72 43.20 0.17 0.23 0.26
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
4 ${AD}_{r}\uparrow$ 99.96 99.99 99.99 99.99 99.99 99.99 80.20 94.82 97.03 22.03 42.26 43.25 0.17 0.25 0.25
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
10 ${AD}_{r}\uparrow$ 99.96 99.99 99.99 99.99 99.99 99.99 75.36 98.07 98.96 19.10 42.60 41.36 0.16 0.27 0.27
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
Table 5. Distributed Data-Parallel Unlearning Performance On CIFAR-100
Model #$C_{f}$ Metrics Original Model (%) Re-training Model (%) FIUn (%) Cumulative Unlearning Time (s)
Re-training FIUn
$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
AlexNet 1 ${AD}_{r}\uparrow$ 98.01 99.96 99.96 99.96 99.96 99.96 97.23 98.85 97.66 30.13 31.26 31.14 0.09 0.14 0.23
${AD}_{f}\downarrow$ 99.99 99.96 99.96 0.00 0.00 0.00 0.00 0.00 0.00
2 ${AD}_{r}\uparrow$ 99.96 99.96 99.96 99.96 99.96 99.96 96.14 97.07 96.71 26.43 26.47 26.41 0.11 0.16 0.27
${AD}_{f}\downarrow$ 99.96 99.96 99.96 0.00 0.00 0.00 0.00 0.00 0.00
4 ${AD}_{r}\uparrow$ 98.01 99.96 99.96 99.96 99.96 99.96 94.58 90.41 90.97 24.33 24.36 24.37 0.10 0.14 0.23
${AD}_{f}\downarrow$ 99.09 99.96 99.96 0.00 0.00 0.00 0.00 0.00 0.00
10 ${AD}_{r}\uparrow$ 98.01 99.96 99.96 99.96 99.96 99.96 91.53 76.19 80.70 21.03 22.96 22.10 0.10 0.14 0.24
${AD}_{f}\downarrow$ 97.15 99.96 99.96 0.00 0.00 0.00 9.73 0.00 11.08
ResNet18 1 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 91.87 94.67 93.67 32.53 31.95 32.18 0.30 0.34 0.64
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
2 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 89.94 94.58 92.40 28.28 27.22 28.47 0.30 0.34 0.66
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
4 ${AD}_{r}\uparrow$ 99.96 89.20 88.52 99.96 99.96 86.44 93.95 89.26 91.77 25.37 24.97 25.13 0.31 0.35 0.66
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
10 ${AD}_{r}\uparrow$ 99.96 89.20 88.52 99.96 99.96 99.96 90.44 95.79 94.19 22.60 21.92 24.19 0.34 0.34 0.68
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 9.73 0.00 0.00
Table 6. Federated Unlearning Performance On TinyImageNet With Multiple Label Distribution
Model #flimit-$\#\cap_{f}$ Metrics Original Model (%) Re-training Model (%) FIUn (%) Cumulative Unlearning Time (s)
Re-training FIUn
$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
DenseNet161 60% $AD_{r}\uparrow$ 99.99 99.99 99.81 98.80 99.99 99.99 96.39 84.75 77.89 3000.66 1998.46 994.32 4.69 2.39 2.30
${AD}_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00
40% $AD_{r}\uparrow$ 99.50 99.99 99.80 99.51 99.99 99.99 96.64 85.40 85.00 3009.53 1010.45 2011.14 5.12 2.69 2.14
${AD}_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00
20% $AD_{r}\uparrow$ 99.99 99.99 99.96 99.16 99.99 99.99 96.22 81.49 78.83 3022.64 2020.67 1021.34 4.98 2.64 2.59
${AD}_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.00 0.00 0.00
0% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.71 96.01 83.04 79.15 3039.64 2041.32 1046.14 4.96 2.14 2.50
${AD}_{f}\downarrow$ 99.99 99.99 98.82 0.00 0.00 0.00 0.00 0.00 0.00
ResNet18 60% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 95.23 98.54 98.44 349.32 246.43 123.74 3.16 1.65 1.68
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
40% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 96.98 99.44 99.49 355.32 247.64 128.43 2.97 1.46 1.59
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
20% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 94.43 98.86 99.38 354.83 246.43 129.35 2.89 1.54 1.38
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
0% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 97.36 99.34 99.32 351.38 244.34 125.32 3.10 1.69 1.39
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
Table 7. Distributed Data-Parallel Unlearning Performance On CIFAR-100 With Multiple Label Distribution
Model #flimit-$\#\cap_{f}$ Metrics Original Model (%) Re-training Model (%) FIUn (%) Cumulative Unlearning Time (s)
Re-training FIUn
$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
DenseNet161 60% $AD_{r}\uparrow$ 99.99 99.99 99.81 98.80 99.99 99.99 95.18 97.07 99.99 999.41 996.43 997.64 0.15 0.08 0.09
${AD}_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00
40% $AD_{r}\uparrow$ 99.99 99.99 99.80 99.51 99.99 99.99 90.53 97.07 98.82 1004.64 1006.31 1006.14 0.19 0.09 0.11
${AD}_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00
20% $AD_{r}\uparrow$ 99.99 99.99 99.96 99.16 99.99 99.99 95.29 97.07 99.89 1019.43 1020.16 1018.64 0.18 0.07 0.10
${AD}_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.00 0.00 0.00
0% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 87.15 97.07 96.36 1031.43 1029.64 1030.64 0.18 0.10 0.11
${AD}_{f}\downarrow$ 99.99 99.99 98.82 0.00 0.00 0.00 0.00 0.00 0.00
ResNet18 60% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 97.62 88.94 87.87 30.36 33.18 37.31 0.60 0.32 0.29
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
40% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 96.50 89.35 87.89 28.30 29.13 52.33 0.59 0.34 0.30
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
20% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 95.31 88.17 93.06 26.39 23.37 56.61 0.58 0.28 0.30
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
0% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 88.87 88.86 89.16 22.03 24.36 46.42 0.64 0.37 0.32
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
5.3. Label Category Settings
The unlearned label categories are represented as #$\#C_{f}$. The distribution of labels among different models may also vary in different data distributions. For instance, model $w_{a}$ aims to unlearn labels 1, 2, 3, and 4, while model $w_{b}$ intends to unlearn labels 2, 3, 5, and 6. In this case, there is an overlap between the unlearned label categories of model $w_{a}$ and model $w_{b}$, specifically labels 2 and 3. We refer to the degree of overlapping among these unlearned label categories as #flimit-$\#\cap_{f}$.
The unlearning process is applied to individual categories and various combinations of labels. For the CIFAR-100 dataset, the combinations we select include 1, 2, 4, and 10 categories. For the TinyImageNet dataset, the selected combinations include 1, 4, 6, and 10 categories. The overlap of the unlearned label categories is divided into different proportions, including 60%, 40%, 20%, and 0%.
6. Evaluation
In this section, we use NVIDIA 4090 GPU to set up a test platform that demonstrates the impact of the FIUn method on the accuracy of unlearned and retained labels in four frameworks: FL, DDPL, IL and TL. When assessing accuracy and time consumption across these learning frameworks during the model unlearning process, we compare our proposed FIUn method with existing unlearning methods and use re-training as a benchmark. We also analyze the impact of the inheritance depth and parameter layers of the models on unlearning effectiveness.
(a) FL in CIFAR100 ResNet18
(b) FL CIFAR100 by AlexNet
(c) FL TinyImageNet by ResNet18
(d) FL TinyImageNet by DenseNet161
Figure 7. Overlap #$\cap_{f}$ analysis, #$C_{f}$=15.
6.1. General Label Unlearning Analysis
We consider existing emerging unlearning methods for comparison: 1) Fine-tuning, which adjusts the original model using the remaining dataset (Odusami et al., 2021), 2) Gradient ascent (GA), employing negative gradients for unlearning (Zhang et al., 2024), and 3) Distill, which distills knowledge from the original model into a student model using the remaining data (Sinha et al., 2023).
Performance comparisons are depicted in Figure 5 for unlearning accuracy $\bigtriangleup_{acc}$ and detailed in Table 1 for unlearning speed. The results demonstrate that our method not only enhances unlearning speed by up to 99% over alternative unlearning methods but also consistently outperforms them in unlearning accuracy.
These results indicate that our method exhibits outstanding effectiveness and efficiency in handling the complexity of model unlearning. In what follows, we evaluate the proposed parallel unlearning method in comparison with the most robust method—re-training. Within all considered learning frameworks, our proposed FIUn method achieves a maximum increase in unlearning speed of 99% compared to the re-training benchmark across different datasets and models. The detailed results are summarized in Tables 2 to 5, with more experiments provided in Appendix .2.1.
1) Single-label unlearning. For all frameworks and model types in CIFAR100, experimental results show that the $AD_{f}$ metric reaches 0, strongly proving that our method effectively unlearns single-class labels. Meanwhile, the average $AD_{r}$ metric is as high as 94.53%, further confirming our method’s efficiency in retaining labels. For all frameworks and model types in TinyImageNet, the $AD_{f}$ metric also reaches 0, and the average $AD_{r}$ metric reaches 79.49%. This significantly validates the effectiveness of our method.
2) Multi-label unlearning. For the CIFAR100 dataset, with #$C_{f}=2,4,10$, the average $AD_{f}$ metric across all frameworks and model types is 1.93%, while the average $AD_{r}$ metric reaches 86.01%. For the TinyImageNet dataset, the average $AD_{f}$ metric is 0.21%, and the average $AD_{r}$ metric is 83.54%. These results further demonstrate the effectiveness of our method. On the other hand, for both the IL and TL frameworks, as the number of unlearning label classes increases, the $AD_{f}$ metric for all models approaches 0. However, the $AD_{r}$ performance of the AlexNet and DenseNet161 models shows a downward trend, while the $AD_{r}$ performance of the ResNet18 model remains stable.
Further observation reveals that the number of parameters in the last layer of DenseNet161 is 4.3 times that of ResNet18, and the number of parameters in the last layer of AlexNet is 7.9 times that of ResNet18. These results indicate that as the number of model parameters and the classes of forgotten labels increases, the $AD_{r}$ performance may decline. This further suggests that while our proposed method effectively unlearns specified labels, it can somewhat affect the accuracy of retained labels.
(a) FL TinyImageNet by ResNet18
(b) FL CIFAR100 by ResNet18
(c) FL TinyImageNet by DenseNet161
(d) FL CIFAR100 by DenseNet161
(e) IL TinyImageNet by ResNet18
(f) IL CIFAR100 by ResNet18
(g) IL TinyImageNet by DenseNet161
(h) IL CIFAR100 by DenseNet161
Figure 8. Inheritance depth analysis #$C_{f}=4$.
(a) FL AlexNet in CIFAR100
(b) FL CIFAR100 by ResNet18
(c) IL CIFAR100 by ResNet18
(d) IL CIFAR100 by AlexNet
(e) FL TinyImageNet by DenseNet161
(f) FL TinyImageNet by ResNet18
(g) IL TinyImageNet by ResNet
(h) IL TinyImageNet by DenseNet161
Figure 9. Unlearning layer analysis #$C_{f}$=15.
6.2. Merged Label Unlearning Analysis
This section assesses the applicability of the merging FIM function to models with various distributions of unlearned labels. We analyze the unlearning performance of different types of models across various datasets within the FL and DDPL frameworks. As demonstrated in Tables 6 and 7, with more experiments provided in Appendix .2.2, the average $AD_{r}$ is 89.72%, and the average $AD_{f}$ is 5.62%, confirming the unlearning effectiveness of our method. We further validate that in the FIM, the impact of data on model parameters remains relatively stable.
We also analyze whether our proposed merging FIM function can handle the merged label unlearning task within the FL framework. As illustrated in Figure 7, in the AlexNet and DenseNet161 models, our method shows an average performance improvement of 88.54%, compared to using FIM directly. In the ResNet18 model, our method’s performance closely matches that of FIM. Additionally, our proposed merged label unlearning is, on average, 2.3 times faster than general label unlearning. This further validates the effiency of our method.
6.3. Inherited Depth of Model Analysis
We further examine whether the performance of FIUn is influenced by the increasing depth of inheritance within the FL and IL frameworks. All experiments are conducted with the same hyperparameters.
  • Federated Learning. In the FL framework, the CIFAR100 and TinyImageNet datasets are randomly divided into five segments. Each model trains using one segment of the data, adopting a binary tree structure as depicted in Figure 6. We evaluate whether our method can effectively unlearn specified labels as the depth of the binary tree increases.
  • Incremental Learning. In the IL framework, as demonstrated in Figure 6, the starting model $w_{g}$ contains data from 170 classes. With each incremental learning step, one class of label data is added to assess whether our method can effectively unlearn specified labels.
As shown in Figure 8 (i.e., single-class unlearning, see Appendix .4), we observe that under both frameworks, various datasets, and models with different unlearned labels, the performance of the proposed algorithm remains consistent regardless of the increased inherited depth. This consistency across both frameworks and models supports the notion that the impact of data on model parameters remains stable at fixed positions. Such stability ensures that our method maintains high performance even as the depth increases.
6.4. Analysis of Parameter Layer Selection in Models
We analyze how the number of model layers used to calculate the FIM impacts the performance of our algorithm. We conduct experiments within both FL and IL frameworks, with 15 label categories selected as unlearned labels. The DAG diagrams for these frameworks are illustrated in Figures 6 and 6. We calculate the FIM using the last layer, the last four layers, the last seven layers, and all layers to assess unlearning performance.
As depicted in Figure 9, we observe that an increase in the number of layers used leads to diminishing unlearning performance across various datasets and models. This decline is attributable to the increase in the layer parameters, which inversely affects the $AD_{r}$ and $AD_{f}$ metrics. Moreover, a larger number of layers prolongs the time needed to compute the FIM. This finding underscores the effectiveness of our method, which relies on calculations from only the last layer.
7. Related Work
Although FL (Li et al., 2020b), DDPL (Li et al., 2020c), IL (Van de Ven et al., 2022), and TL (Weiss et al., 2016) have each made significant strides in addressing specific challenges, such as data privacy protection (Li et al., 2021), model training efficiency (Dean et al., 2012), knowledge retention and unlearning (Han et al., 2023), and cross-domain knowledge transfer (Rajapaksha et al., 2023), systematic exploration that combines these learning frameworks remains sparse. Most studies have merely attempted simple combinations of a few methods (Dai and Meng, 2023; Yoon et al., 2021; Dai et al., 2019; Miao et al., 2023) without touching the core of their commonality–model inheritance. This characteristic plays a crucial role in environments where unlearning is conducted on a large-scale. As data and models grow, the computational overhead increases exponentially with the depth of inheritance, while time consumption grows linearly with the depth, even when unlearning can be executed simultaneously at the same depth starting from the unlearning origin. This holds true regardless of the existing unlearning methods utilized, which are respectively discussed in the following.
Federated Unlearning. FedEraser (Liu et al., 2020) used historical parameter updates from the central server to reconstruct the deleted model. Wu et al. (Wu et al., 2022a) subtracted historical updates and used knowledge distillation to restore model performance without client data. FRU (Yuan et al., 2023) recalled and calibrated historical updates to eliminate user contributions and accelerate federated recommendation reconstruction. These methods primarily relied on recalling and calibrating historical parameter updates of the model to achieve deletion or re-training. Halimi et al. (Halimi et al., 2022) reversed the learning process by maximizing local empirical loss and used PGD to perform deletion on the client side, while Wu et al. (Wu et al., 2022b) used class, client, and sample decoupling learning combined with Stochastic Gradient Ascent (SGA) and Elastic Weight Consolidation (EWC) to achieve deletion. These methods achieved deletion by optimizing the model’s loss function or gradients.
Liu et al. (Liu et al., 2022) utilized first-order Taylor expansion and a diagonal empirical Fisher information matrix for accelerated re-training. Wang et al. (Wang et al., 2022) applied CNN channel pruning to remove specific category information, while FFMU (Che et al., 2023) employed nonlinear function analysis to enhance model deletion processes. These approaches streamline deletion or re-training by altering the model structure or employing targeted techniques. Additionally, KNOT (Su and Li, 2023) optimized re-training by introducing clustering aggregation and framing the client clustering issue as a dictionary minimization problem.
Distributed Data-Parallel Unlearning. In DDPL knowledge unlearning, handling the unlearning process is restricted to the data shard on the specific device, followed by model aggregation and iterative training. If unlearning is required from a specific shard, only the model trained on that shard is re-trained. Approximate machine unlearning directly modifies trained model parameters, employing methods such as 1) Gradient-based approaches (Fan et al., 2023; Golatkar et al., 2020; Neel et al., 2021), which update model gradients to gradually diminish the data’s influence, and 2) Data perturbation (Tarun et al., 2023; Goel et al., 2024), which introduces noise to reduce specific data points’ impact on the model. These methods facilitate effective knowledge unlearning in DDPL.
Incremental Unlearning. Research on unlearning on IL is still nascent, yet the Projected-Gradient Unlearning method (PGU) (Hoang et al., 2024) shows promise. PGU computes the gradients for the unlearning dataset and projects these onto the Core Gradient Space (CGS) derived from the training set. It then subtracts this projected component from the original gradients, positioning the residuals in a space orthogonal to the CGS, and updates the model parameters with these adjusted gradients. Effective for one-time data deletions and batch-wise unlearning, PGU is well-suited for ongoing incremental learning processes.
Transfer Unlearning. Research on unlearning specific target task knowledge has been relatively scarce in TL. To address this challenge, (Sepahvand et al., 2024) proposed fine-tuning pre-trained models through data selection. Subsequently, the entire network was fine-tuned on the target task using gradient descent.
What Sets Our Solution Apart. The above-mentioned existing studies on unlearning methods typically endure high computational complexity, significant resource demands, and implementation difficulties. These methods are time-consuming and resource-intensive, especially when deployed in large-scale environments, as they require accurate tracking and sequential updating of numerous model parameters. By contrast, the proposed FIUn method fully parallelizes unlearning, significantly enhancing computational efficiency, resource utilization, and adaptability. It effectively addresses the challenges faced by existing unlearning methods in inheritance-oriented model networks.
8. Conclusion
In this paper, we developed a new parallel unlearning framework for models with dependence and inheritance. To do this, we introduced the UMIG, which can effectively abstract model inheritance, e.g., in FL, DDPL, IL, and TL, using a DAG topology. We also developed the FIUn method and the merging FIM function, which can quickly locate impacted models and conduct parallel unlearning through the UMIG. Experimental results on various models and datasets showed that our method achieves complete unlearning for single-class labels, while maintaining an average accuracy of 94.53% for retained labels. For multi-class labels, the unlearning accuracy is only 1.07%, with retained label accuracy at 84.77%. Our method is 99% faster than existing unlearning methods, efficiently handling model updates regardless of inheritance depth.